OpenStudy (natalie1234):

Will give medal and fan!! Find a number that is between 6/11 and 0.67 A. 58/99 B. 69/99 C. 24/33 D. 26/55

OpenStudy (anonymous):

\[A. 58/99 = .58585858\]\[B. 69/99 = .696969\]\[c. 24/33 = .7272727\]\[D. 26/55 = .47272727\]\[6/11= .54 < x < .67\]

OpenStudy (natalie1234):

so is it a?

OpenStudy (anonymous):

Yup

OpenStudy (natalie1234):

thanks (:

OpenStudy (anonymous):

NP :)